# Intrinsic-Metric Physics-Informed Neural Networks (IM-PINN) for Reaction-Diffusion Dynamics on Complex Riemannian Manifolds

### v1

@@ Abstract
Simulating nonlinear reaction-diffusion dynamics on complex, non-Euclidean manifolds remains a fundamental challenge in computational morphogenesis, constrained by high-fidelity mesh generation costs and symplectic drift in discrete time-stepping schemes. This study introduces the Intrinsic-Metric Physics-Informed Neural Network (IM-PINN), a mesh-free geometric deep learning framework that solves partial differential equations directly in the continuous parametric domain. By embedding the Riemannian metric tensor into the automatic differentiation graph, our architecture analytically reconstructs the Laplace-Beltrami operator, decoupling solution complexity from geometric discretization. We validate the framework on a "Stochastic Cloth" manifold with extreme Gaussian curvature fluctuations ($K\in[-2489,3580]$), where traditional adaptive refinement fails to resolve anisotropic Turing instabilities. Using a dual-stream architecture with Fourier feature embeddings to mitigate spectral bias, the IM-PINN recovers the "splitting spot" and "labyrinthine" regimes of the Gray-Scott model. Benchmarking against the Surface Finite Element Method (SFEM) reveals superior physical rigor: the IM-PINN achieves global mass conservation error of $\mathcal{E}_{mass}\approx 0.157$ versus SFEM’s $0.258$, acting as a thermodynamically consistent global solver that eliminates mass drift inherent in semi-implicit integration. The framework offers a memory-efficient, resolution-independent paradigm for simulating biological pattern formation on evolving surfaces, bridging differential geometry and physics-informed machine learning.
@@ 1 Introduction
The remainder of this manuscript is meticulously structured to establish both the theoretical rigor and the computational utility of the proposed framework. Section 2 (Methodology) provides the foundational mathematical derivation of the "Stochastic Cloth," explicitly defining the parametric height functions that induce multi-scale Gaussian perturbations. Here, we formulate the reaction-diffusion dynamics via intrinsic differential operators, detailing the construction of the Riemannian metric tensor $g_{ij}$ and the corresponding "Dual-Stream" neural architecture equipped with Fourier feature embeddings and the mass-conservative loss function. Section 3 (Results) presents a comprehensive empirical evaluation, benchmarking the IM-PINN against the Surface Finite Element Method (SFEM) [9] across a battery of quantitative metrics, including pointwise tracking error, manifold capacity, and pattern topology verification. Section 4 (Discussion) synthesizes these findings, offering a phenomenological analysis of how intrinsic curvature modulates Turing instability and addressing the stability concerns raised in recent physics-informed learning literature [34] through the lens of thermodynamic consistency. Finally, Section 5 (Conclusion) summarizes the transformative potential of mesh-free geometric deep learning and outlines strategic avenues for future research, particularly regarding time-evolving domains and inverse shape design.
@@ 2.4 High-Fidelity Numerical Baseline and Error Metrics
In the absence of closed-form analytical solutions for nonlinear reaction-diffusion systems on arbitrary Riemannian manifolds, establishing a rigorous ground truth is paramount for validating the proposed learning framework. To this end, we generate a high-fidelity reference solution using the Surface Finite Element Method (SFEM), widely regarded as the gold standard for geometric PDEs [9]. We discretize the continuous stochastic manifold $\mathcal{M}$ into a high-resolution triangular mesh $\mathcal{T}_{h}$ comprising over 40,000 vertices to ensure mesh independence and capture the finest scales of the Gaussian perturbations (see Figure 1). The governing equations are cast into their weak variational formulation, where the solution is sought in the Sobolev space $H^{1}(\mathcal{M})$ via a semi-implicit Euler time-stepping scheme to handle the stiffness of the reaction terms [34]. Specifically, the diffusion operator is approximated using the cotangent stiffness matrix, while the nonlinear reaction kinetics are linearized locally to ensure unconditional stability. The resulting discrete solution vectors $U_{ref}$ and $V_{ref}$ serve as the ground truth for benchmarking. To quantitatively assess the performance of the IM-PINN, we compute the relative $\mathcal{L}_{2}$ error norm integrated over the manifold surface:
@@ 2.5 Spatially Modulated Constitutive Relations and Parameter Mapping
(a) Intrinsic Potential $\phi(u,v)$
(b) Manifold Potential $\phi(\mathbf{r}(u,v))$
Figure 2: Spatial Modulation of the Chemical Potential. (a) The scalar field $\phi(u,v)$ defined in the parametric domain $\Omega=[0,1]^{2}$, showing the intrinsic periodic structure before geometric deformation. (b) The same field mapped onto the stochastic cloth manifold $\mathcal{M}$. The color intensity represents the magnitude of the potential, which locally modulates the feed rate $F$ of the Gray-Scott system. This visualization confirms that the IM-PINN must learn to resolve pattern variations that are driven by both the extrinsic curvature of the surface (wrinkles) and this intrinsic chemical heterogeneity.
@@ 2.8 Algorithmic Implementation and Training Protocol
Algorithm 1 Training the Intrinsic-Metric PINN (IM-PINN)
@@ 2.9 Architectural Schema and Computational Graph
The architecture is fundamentally bifurcated into two parallel processing streams that converge only at the loss evaluation stage. The Neural Stream (Top) is responsible for function approximation. It ingests the raw spatiotemporal coordinates $(u,v,t)$ and immediately projects them into a high-dimensional Fourier Feature space via the mapping $\gamma(\mathbf{x})$. This step is mathematically critical; by lifting the low-dimensional inputs onto a hypersphere of sinusoidal functions, we artificially induce a stationary kernel in the Neural Tangent Kernel (NTK) limit, empowering the subsequent Multi-Layer Perceptron (MLP) to resolve the high-frequency spectral components of the Turing instabilities that would otherwise be smoothed out by spectral bias.
@@
The critical innovation lies in the convergence of these streams at the Automatic Differentiation (AD) Module. Here, the "dumb" gradients of the neural network outputs ($\nabla\hat{U},\nabla\hat{V}$)—which are computed with respect to the flat parametric space—are "geometry-corrected" via contraction with the inverse metric tensor from the geometric stream. This synthesis yields the true Laplace-Beltrami operator $\Delta_{\mathcal{M}}$, ensuring that the diffusion physics propagates along the geodesic paths of the curved surface rather than the Euclidean shortcuts of the embedding space. Furthermore, the diagram explicitly visualizes the injection of the Chemical Potential field $\phi$, which enters the graph as an external modulator, spatially distorting the reaction kinetics term $R(U,V)$ within the PDE residuals. This end-to-end differentiable pipeline ensures that the backpropagated error signal $\nabla_{\theta}\mathcal{L}$ carries precise geometric information, forcing the network weights $\theta$ to conform to the intrinsic topology of the manifold.
@@ 3.1 Turing Pattern Topology and Geometric Modulation
Table 2: Geometric Statistics of the Stochastic Cloth Manifold. The extreme range of the Gaussian Curvature ($K$) and the Mean Curvature ($H$) demonstrates the challenge posed by the surface. The high standard deviation of the metric determinant $\det(g)$ indicates significant local area distortion that the IM-PINN must normalize via the $\sqrt{|g|}$ term.
@@
(a) Species $U$ (Inhibitor)
(b) Species $V$ (Activator)
@@ 3.2 Quantitative Fidelity and Conservation Metrics
Table 3: Quantitative Comparison: IM-PINN vs. SFEM Baseline. The metrics highlight the trade-off inherent in neural physics solvers. While the pointwise $\mathcal{L}_{2}$ error grows due to the chaotic phase sensitivity of the Turing patterns, the IM-PINN demonstrates superior conservation of physical invariants (lower Mass Error) and significantly reduced memory complexity compared to the dense matrix operations of the SFEM.
@@ 3.3 Convergence Dynamics and Computational Efficiency
Figure 5: Training Dynamics of the IM-PINN. (a) The total loss shows rapid initial convergence followed by fine-tuning. (b) The PDE residual loss drops by seven orders of magnitude, confirming that the network satisfies the local governing equations. (d) The component breakdown reveals that the boundary conditions ($\mathcal{L}_{BC}$, green) are satisfied almost to machine precision ($\sim 10^{-8}$), while the Mass Loss ($\mathcal{L}_{Mass}$, red) acts as a persistent regularizer, preventing the solution from drifting into unphysical regimes.
@@ 3.4 Spatial Error Distribution and Geometric Correlation
(a) Predicted Species $U$
(b) Absolute Error $|U_{pinn}-U_{sfem}|$
(c) Predicted Species $V$
@@ 3.5 Mechanistic Analysis of Flux Dynamics and Reaction Kinetics
(a) Diffusion Flux $|\nabla_{\mathcal{M}}U|$
(b) Reaction Rate $\mathcal{R}(U,V)$
(c) Chemical Potential $\phi(\xi)$
@@ 3.6 Comparative Computational Complexity and Resolution Independence
Finally, we evaluate the scalability of the IM-PINN framework relative to the spectral-geometric baseline. A fundamental limitation of the Surface Finite Element Method (SFEM) is its rigid coupling between spatial resolution and computational cost; refining the mesh to capture microscopic Gaussian perturbations ($\sigma\ll 1$) necessitates a quadratic increase in the stiffness matrix size, leading to prohibitive memory requirements for high-frequency manifolds. In stark contrast, the IM-PINN learns a continuous functional representation of the solution $\hat{U}(\xi,t)$, decoupled from any discrete tessellation. As detailed in the comprehensive performance summary (Table 4), the neural solver maintains a fixed memory footprint of 82,690 parameters (approx. 330 KB) regardless of the geometric complexity, whereas the SFEM baseline requires storing mesh connectivity and quadrature data scaling linearly with the vertex count ($N_{v}>40,000$). This "resolution independence" allows the trained IM-PINN to act as a super-resolution operator, enabling the inference of chemical concentrations at arbitrary coordinate locations without re-training—a capability we define as infinite-resolution query. While the offline training latency of the IM-PINN (773s) exceeds the forward integration time of the SFEM (0.54s/step), this initial cost is amortized by the model’s ability to enforce stricter conservation laws ($\mathcal{E}_{mass}^{PINN}<\mathcal{E}_{mass}^{SFEM}$) and its seamless differentiability for potential inverse design tasks. Consequently, the IM-PINN represents a paradigm shift from "compute-bound" simulation to "memory-efficient" neural representation, particularly advantageous for storage-constrained environments or multi-scale geometries where meshing is intractable.
@@ 4 Discussion
The successful deployment of the Intrinsic-Metric Physics-Informed Neural Network (IM-PINN) on the highly irregular "Stochastic Cloth" manifold represents a significant departure from the discretization-heavy paradigms that have dominated computational morphogenesis for decades. Historically, the numerical treatment of Reaction-Diffusion (RD) systems on curved surfaces has been inextricably linked to the quality of the underlying mesh triangulation [9]. In traditional frameworks like the Surface Finite Element Method (SFEM) or Discrete Exterior Calculus (DEC), the geometric fidelity is capped by the polynomial order of the basis functions and the aspect ratio of the elements; highly curved regions—such as the Gaussian wrinkles explored in this study—often necessitate adaptive mesh refinement (AMR) to prevent numerical locking or artificial dissipation [3]. Our results demonstrate that the IM-PINN circumvents these topological constraints entirely by learning a mesh-free, continuous functional representation of the state variables. By encoding the Riemannian metric $g_{ij}$ directly into the automatic differentiation graph, we effectively replace the discrete stiffness matrix with an exact, analytical evaluation of the Laplace-Beltrami operator. This methodological shift aligns with the emerging principles of "Geometric Deep Learning," where the inductive bias of the network is tailored to respect the invariance groups of the non-Euclidean domain [10]. Furthermore, our use of Fourier Feature Embeddings to mitigate the spectral bias of the Neural Tangent Kernel (NTK) [30] proved decisive in resolving the sharp wavefronts of the Turing patterns. Where standard MLPs fail to capture high-frequency spatial oscillations, our spectrally enriched architecture successfully recovered the anisotropic "splitting spots" regime (Figure 4), validating the hypothesis that coordinate-based neural networks can function as universal approximators for geometric PDEs provided the input space is appropriately lifted [16].
@@ 5 Conclusion
This research has successfully demonstrated that the Intrinsic-Metric Physics-Informed Neural Network (IM-PINN) constitutes a viable and rigorously physical paradigm for simulating complex reaction-diffusion dynamics on manifolds with extreme geometric irregularity. By coupling the spectral properties of Fourier feature embeddings with the differential geometry of the Riemannian metric tensor, we have overcome the two primary bottlenecks of classical computational morphogenesis: the "spectral bias" that smooths out high-frequency Turing patterns, and the "meshing bottleneck" that mandates prohibitive discretization densities in regions of high curvature. The successful emergence of anisotropic "splitting spots" on the "Stochastic Cloth"—a surface exhibiting Gaussian curvature fluctuations spanning four orders of magnitude ($K\in[-2489,3580]$)—validates our core hypothesis: that the coordinate-based neural network can serve not merely as an interpolator, but as a fully differentiable, resolution-independent operator that solves the governing PDEs directly in the continuous parametric domain. This achievement effectively decouples the complexity of the solution field from the complexity of the domain geometry, offering a scalable alternative to the Surface Finite Element Method (SFEM) for high-frequency manifolds. A definitive finding of this study is the superior thermodynamic consistency of the neural solver compared to traditional discrete variational methods. While the chaotic nature of the Gray-Scott system renders long-term pointwise tracking notoriously difficult—evidenced by the inevitable phase decoherence between the IM-PINN and the SFEM ground truth—the global conservation metrics reveal a critical advantage. The IM-PINN achieved a mass violation error of $\mathcal{E}_{mass}\approx 0.157$, significantly outperforming the baseline error of $0.258$. This result overturns the prevailing skepticism that neural networks are "unphysical black boxes." We argue that by treating the mass balance equation not as a post-hoc check but as a primary component of the loss landscape ($\mathcal{L}_{Mass}$), the IM-PINN acts as a "global solver," enforcing conservation laws simultaneously across the entire spatiotemporal volume. This contrasts sharply with semi-implicit time-stepping schemes, which are prone to the monotonic accumulation of truncation errors (mass drift) over long integration horizons. Thus, for biological applications where the strict conservation of morphogen quantity is more critical than the precise location of every microscopic spot, the IM-PINN offers a more physically robust framework.

### v2

@@ Abstract
Simulating nonlinear reaction-diffusion dynamics on complex, non-Euclidean manifolds remains a fundamental challenge in computational morphogenesis, constrained by high-fidelity mesh generation costs and numerical drift in discrete time-stepping schemes. This study introduces the Intrinsic-Metric Physics-Informed Neural Network (IM-PINN), a mesh-free geometric deep learning framework that solves partial differential equations directly in the continuous parametric domain. By embedding the Riemannian metric tensor into the automatic differentiation graph, our architecture analytically computes the Laplace-Beltrami operator, decoupling solution complexity from geometric discretization. We validate the framework on a "Stochastic Cloth" manifold with extreme Gaussian curvature fluctuations ($K\in[-2489,3580]$), where traditional adaptive refinement fails to resolve anisotropic Turing instabilities. Using a dual-stream architecture with Fourier feature embeddings to mitigate spectral bias, the IM-PINN recovers the "splitting spot" and "labyrinthine" regimes of the Gray-Scott model. Benchmarking against the Surface Finite Element Method (SFEM) reveals superior physical rigor: the IM-PINN achieves global mass conservation error of $\mathcal{E}_{mass}\approx 0.157$ versus SFEM’s $0.258$, acting as a thermodynamically consistent global solver that eliminates mass drift inherent in semi-implicit integration. The framework offers a memory-efficient, resolution-independent paradigm for simulating biological pattern formation on evolving surfaces, bridging differential geometry and physics-informed machine learning.
@@ 1 Introduction
The remainder of this manuscript is meticulously structured to establish both the theoretical rigor and the computational utility of the proposed framework. Section 2 (Methodology) provides the foundational mathematical derivation of the "Stochastic Cloth," explicitly defining the parametric height functions that induce multi-scale Gaussian perturbations. Here, we formulate the reaction-diffusion dynamics via intrinsic differential operators, detailing the construction of the Riemannian metric tensor $g_{ij}$ and the corresponding "Dual-Stream" neural architecture equipped with Fourier feature embeddings and the mass-conservative loss function. Section 3 (Results) presents a comprehensive empirical evaluation, benchmarking the IM-PINN against the SFEM [9] across a battery of quantitative metrics, including pointwise tracking error, manifold capacity, and pattern topology verification. Section 4 (Discussion) synthesizes these findings, offering a phenomenological analysis of how intrinsic curvature modulates Turing instability and addressing the stability concerns raised in recent physics-informed learning literature [31] through the lens of thermodynamic consistency. Finally, Section 5 (Conclusion) summarizes the transformative potential of mesh-free geometric deep learning and outlines strategic avenues for future research, particularly regarding time-evolving domains and inverse shape design.
@@ 2.4 High-Fidelity Numerical Baseline and Error Metrics
In the absence of closed-form analytical solutions for nonlinear reaction-diffusion systems on arbitrary Riemannian manifolds, establishing a rigorous ground truth is paramount for validating the proposed learning framework. To this end, we generate a high-fidelity reference solution using the SFEM, widely regarded as the gold standard for geometric PDEs [9]. We discretize the continuous stochastic manifold $\mathcal{M}$ into a high-resolution triangular mesh $\mathcal{T}_{h}$ comprising over 40,000 vertices to ensure mesh independence and capture the finest scales of the Gaussian perturbations (see Figure 1). The governing equations are cast into their weak variational formulation, where the solution is sought in the Sobolev space $H^{1}(\mathcal{M})$ via a semi-implicit Euler time-stepping scheme to handle the stiffness of the reaction terms [31]. Specifically, the diffusion operator is approximated using the cotangent stiffness matrix, while the nonlinear reaction kinetics are linearized locally to ensure unconditional stability. The resulting discrete solution vectors $U_{ref}$ and $V_{ref}$ serve as the ground truth for benchmarking. To quantitatively assess the performance of the IM-PINN, we compute the relative $\mathcal{L}_{2}$ error norm integrated over the manifold surface:
@@ 2.5 Spatially Modulated Constitutive Relations and Parameter Mapping
(a) Intrinsic Potential $\phi(u,v)$, defined as $\phi(u,v)=\cos(4\pi u)\sin(4\pi v)$, represents a spatially varying constitutive parameter.
(b) Manifold Potential $\phi(\mathbf{r}(u,v))$: Represents the scalar chemical potential field $\phi:\Omega\to[-1,1]$, defined as $\phi(u,v)=\cos(4\pi u)\sin(4\pi v)$.
Figure 2: Spatial Modulation of the Chemical Potential. (a) The scalar field $\phi(u,v)$ defined in the parametric domain $\Omega=[0,1]^{2}$, showing the intrinsic periodic structure before geometric deformation. (b) The same field mapped onto the stochastic cloth manifold $\mathcal{M}$. The color intensity represents the magnitude of the potential, which locally modulates the feed rate $F$ of the Gray-Scott system. This visualization confirms that the IM-PINN must learn to resolve pattern variations that are driven by both the extrinsic curvature of the surface (wrinkles) and this intrinsic chemical heterogeneity, as defined by $\phi(u,v)=\cos(4\pi u)\sin(4\pi v)$ in the parametric domain.
@@ 2.8 Algorithmic Implementation and Training Protocol
Algorithm 1 Training the Intrinsic-Metric PINN (IM-PINN) framework
@@ 2.9 Architectural Schema and Computational Graph
The architecture is fundamentally bifurcated into two parallel processing streams that converge only at the loss evaluation stage. The Neural Stream (Top) is responsible for function approximation. It ingests the raw spatiotemporal coordinates $(u,v,t)$ and immediately projects them into a high-dimensional Fourier Feature space via the mapping $\gamma(\mathbf{x})$. This step is mathematically critical; by lifting the low-dimensional inputs onto a hypersphere of sinusoidal functions, we artificially induce a stationary kernel in the NTK limit, empowering the subsequent MLP to resolve the high-frequency spectral components of the Turing instabilities that would otherwise be smoothed out by spectral bias.
@@
The critical innovation lies in the convergence of these streams at the AD Module. Here, the "dumb" gradients of the neural network outputs ($\nabla\hat{U},\nabla\hat{V}$)—which are computed with respect to the flat parametric space—are "geometry-corrected" via contraction with the inverse metric tensor from the geometric stream. This synthesis yields the true Laplace-Beltrami operator $\Delta_{\mathcal{M}}$, ensuring that the diffusion physics propagates along the geodesic paths of the curved surface rather than the Euclidean shortcuts of the embedding space. Furthermore, the diagram explicitly visualizes the injection of the Chemical Potential field $\phi$, which enters the graph as an external modulator, spatially distorting the reaction kinetics term $R(U,V)$ within the PDE residuals. This end-to-end differentiable pipeline ensures that the backpropagated error signal $\nabla_{\theta}\mathcal{L}$ carries precise geometric information, forcing the network weights $\theta$ to conform to the intrinsic topology of the manifold.
@@ 3.1 Turing Pattern Topology and Geometric Modulation
Table 2: Geometric Statistics of the Stochastic Cloth Manifold. The extreme range of the Gaussian Curvature ($K$, spanning from 2489-2489 to +3580+3580 with $\mu_{K}=-3.75$, $\sigma_{K}=447.7$) and the Mean Curvature ($H$) demonstrates the challenge posed by the surface. The high standard deviation of the metric determinant $\det(g)$ indicates significant local area distortion that the IM-PINN must normalize via the $\sqrt{|g|}$ term. The extreme range of the Gaussian Curvature ($K$) and the Mean Curvature ($H$) demonstrates the challenge posed by the surface. The high standard deviation of the metric determinant $\det(g)$ indicates significant local area distortion that the IM-PINN must normalize via the $\sqrt{|g|}$ term.
@@
(a) Species $U$ (Inhibitor) concentration field at $t=2000$.
(b) Species $V$ (Activator) at $t=2000$
@@ 3.2 Quantitative Fidelity and Conservation Metrics
Table 3: Quantitative Comparison: IM-PINN vs. SFEM Baseline. The metrics highlight the trade-off inherent in neural physics solvers. While the pointwise $\mathcal{L}_{2}$ error grows due to the chaotic phase sensitivity of the Turing patterns, the IM-PINN demonstrates superior conservation of physical invariants (lower Mass Error, $\mathcal{E}_{mass}=0.157$) and significantly reduced memory complexity compared to the dense matrix operations of the SFEM (Mass Error, $\mathcal{E}_{mass}=0.258$).
@@ 3.3 Convergence Dynamics and Computational Efficiency
Figure 5: Training Dynamics of the IM-PINN. (a) The total loss shows rapid initial convergence followed by fine-tuning. (b) The PDE residual loss drops by seven orders of magnitude, confirming that the network satisfies the local governing equations. (d) The component breakdown reveals that the boundary conditions ($\mathcal{L}_{BC}$, green) are satisfied almost to machine precision ($\sim 10^{-8}$), the Mass Loss ($\mathcal{L}_{Mass}$, red) stabilizes at $\sim 0.22$ and acts as a persistent regularizer, preventing the solution from drifting into unphysical regimes, and the PDE residual ($\mathcal{L}_{PDE}$) converges to $\sim 3\times 10^{-4}$.
@@ 3.4 Spatial Error Distribution and Geometric Correlation
(a) Predicted Species $U$ at $t=2000$
(b) Absolute Error $|U_{pinn}-U_{sfem}|$ (concentration units)
(c) Predicted steady-state concentration fields for Species $V$ at $t=2000$
@@ 3.5 Mechanistic Analysis of Flux Dynamics and Reaction Kinetics
(a) Diffusion Flux $|abla_{\mathcal{M}}U|=\sqrt{g^{ij}\partial_{i}U\partial_{j}U}$.
(b) Reaction Rate $\mathcal{R}(U,V)$ (inhibitor, activator)
(c) Chemical Potential $\phi(\xi)$ in the parametric domain $\Omega=[0,1]^{2}$.
@@ 3.6 Comparative Computational Complexity and Resolution Independence
Finally, we evaluate the scalability of the IM-PINN framework relative to the spectral-geometric baseline. A fundamental limitation of the SFEM is its rigid coupling between spatial resolution and computational cost; refining the mesh to capture microscopic Gaussian perturbations ($\sigma\ll 1$) necessitates a quadratic increase in the stiffness matrix size, leading to prohibitive memory requirements for high-frequency manifolds. In stark contrast, the IM-PINN learns a continuous functional representation of the solution $\hat{U}(\xi,t)$, decoupled from any discrete tessellation. As detailed in the comprehensive performance summary (Table 4), the neural solver maintains a fixed memory footprint of 82,690 parameters (approx. 330 KB) regardless of the geometric complexity, whereas the SFEM baseline requires storing mesh connectivity and quadrature data scaling linearly with the vertex count ($N_{v}>40,000$). This "resolution independence" allows the trained IM-PINN to act as a super-resolution operator, enabling the inference of chemical concentrations at arbitrary coordinate locations without re-training—a capability we define as infinite-resolution query. While the offline training latency of the IM-PINN (773s) exceeds the forward integration time of the SFEM (0.54s/step), this initial cost is amortized by the model’s ability to enforce stricter conservation laws ($\mathcal{E}_{mass}^{PINN}<\mathcal{E}_{mass}^{SFEM}$) and its seamless differentiability for potential inverse design tasks. Consequently, the IM-PINN represents a paradigm shift from "compute-bound" simulation to "memory-efficient" neural representation, particularly advantageous for storage-constrained environments or multi-scale geometries where meshing is intractable.
@@ 4 Discussion
The successful deployment of the IM-PINN on the highly irregular "Stochastic Cloth" manifold represents a significant departure from the discretization-heavy paradigms that have dominated computational morphogenesis for decades. Historically, the numerical treatment of Reaction-Diffusion (RD) systems on curved surfaces has been inextricably linked to the quality of the underlying mesh triangulation [9]. In traditional frameworks like the SFEM or Discrete Exterior Calculus (DEC), the geometric fidelity is capped by the polynomial order of the basis functions and the aspect ratio of the elements; highly curved regions—such as the Gaussian wrinkles explored in this study—often necessitate adaptive mesh refinement (AMR) to prevent numerical locking or artificial dissipation [3]. Our results demonstrate that the IM-PINN circumvents these topological constraints entirely by learning a mesh-free, continuous functional representation of the state variables. By encoding the Riemannian metric $g_{ij}$ directly into the automatic differentiation graph, we effectively replace the discrete stiffness matrix with an exact, analytical evaluation of the Laplace-Beltrami operator. This methodological shift aligns with the emerging principles of "Geometric Deep Learning," where the inductive bias of the network is tailored to respect the invariance groups of the non-Euclidean domain [10]. Furthermore, our use of Fourier Feature Embeddings to mitigate the spectral bias of the NTK [27] proved decisive in resolving the sharp wavefronts of the Turing patterns. Where standard MLPs fail to capture high-frequency spatial oscillations, our spectrally enriched architecture successfully recovered the anisotropic "splitting spots" regime (Figure 4), validating the hypothesis that coordinate-based neural networks can function as universal approximators for geometric PDEs provided the input space is appropriately lifted [16].
@@ 5 Conclusion
This research has successfully demonstrated that the IM-PINN constitutes a viable and rigorously physical paradigm for simulating complex reaction-diffusion dynamics on manifolds with extreme geometric irregularity. By coupling the spectral properties of Fourier feature embeddings with the differential geometry of the Riemannian metric tensor, we have overcome the two primary bottlenecks of classical computational morphogenesis: the "spectral bias" that smooths out high-frequency Turing patterns, and the "meshing bottleneck" that mandates prohibitive discretization densities in regions of high curvature. The successful emergence of anisotropic "splitting spots" on the "Stochastic Cloth"—a surface exhibiting Gaussian curvature fluctuations spanning four orders of magnitude ($K\in[-2489,3580]$)—validates our core hypothesis: that the coordinate-based neural network can serve not merely as an interpolator, but as a fully differentiable, resolution-independent operator that solves the governing PDEs directly in the continuous parametric domain. This achievement effectively decouples the complexity of the solution field from the complexity of the domain geometry, offering a scalable alternative to the SFEM for high-frequency manifolds. A definitive finding of this study is the superior thermodynamic consistency of the neural solver compared to traditional discrete variational methods. While the chaotic nature of the Gray-Scott system renders long-term pointwise tracking notoriously difficult—evidenced by the inevitable phase decoherence between the IM-PINN and the SFEM ground truth—the global conservation metrics reveal a critical advantage. The IM-PINN achieved a mass violation error of $\mathcal{E}_{mass}\approx 0.157$, significantly outperforming the baseline error of $0.258$. This result overturns the prevailing skepticism that neural networks are "unphysical black boxes." We argue that by treating the mass balance equation not as a post-hoc check but as a primary component of the loss landscape ($\mathcal{L}_{Mass}$), the IM-PINN acts as a "global solver," enforcing conservation laws simultaneously across the entire spatiotemporal volume. This contrasts sharply with semi-implicit time-stepping schemes, which are prone to the monotonic accumulation of truncation errors (mass drift) over long integration horizons. Thus, for biological applications where the strict conservation of morphogen quantity is more critical than the precise location of every microscopic spot, the IM-PINN offers a more physically robust framework.